# Exact diagonalization study of energy level statistics in harmonically confined interacting bosons

### v1

@@ IV.2.2 Long-range correlations
Figure 5: (Color online) Upper panel: Nearest-neighbor spacing distribution $P(s)$; Lower panel: Distribution of the ratio of consecutive level spacings $P(r)$ for strong interaction regime with $g_{2}=7.301$ in the single-vortex state $L_{z}=N$ for $N=12$, 161616, and 202020. The histogram in each graphs represents our numerical result for the lowest 100 energy levels. The blue line shows Poisson, the orange dashed line GOE, and the green dash-dotted line Brody distribution with fitted parameter $b$.
(a) $L_{z}=N$, $g_{2}=0.7301$
(b) $L_{z}=N$, $g_{2}=0.7301$
(c) $L_{z}=N$, $g_{2}=7.301$
(d) $L_{z}=N$, $g_{2}=7.301$
Figure 6: (Color online) Spectral average $\langle\Delta_{3}(L)\rangle$ vs. $L$ and number variance $\Sigma^{2}(L)$ vs. $L$ for moderate interaction with $g_{2}=0.7301$( upper panel) and strong interaction with $g_{2}=7.301$ (lower panel) regimes in the single-vortex state $L_{z}=N$ for $N=12$, 161616, and 202020. Blue circles, green squares, and red diamonds show numerical results. Black dashed and magenta dash-dotted lines correspond to Poisson and GOE results, respectively (The inset in Fig. (a) displays a zoomed-in view up to $L=15$).
@@
Figure 7: (Color online) The distribution of the ratio of consecutive level spacings $P(r)$ for strong interaction regime with $g_{2}=7.301$ for $N=12,16,20$ in $L_{z}=2N$ (Upper Panel) and $L_{z}=3N$ (Lower Panel). The histogram in each graphs represents our numerical result for the lowest 100 energy levels. The blue line corresponds to the Poisson distribution, the orange dashed curve to the GOE distribution.
(a) $L_{z}=2N$, $g_{2}=7.301$
(b) $L_{z}=3N$, $g_{2}=7.301$
Figure 8: (Color online) The spectral average $\langle\Delta_{3}(L)\rangle$ vs $L$ is presented for strong interaction regime with $g_{2}=7.301$ for different number of bosons $N=12,16$ and 202020 in $L_{z}=2N$ (Left) and $L_{z}=3N$ (Right). The blue circle, green square and red diamond lines are our numerical results of $\langle\Delta_{3}(L)\rangle$ for $N=12,16$ and 202020, respectively, for the lowest 100 energy levels . For reference, we have also drawn $\langle\Delta_{3}(L)\rangle$ as a function of $L$, corresponding to the Poisson distribution (black dashed line) and the GOE distribution (magenta dash-dot line).
@@
For small values of energy interval $L$, the numerical data of $\Delta_{3}(L)$ statistic for $N=12,16,20$ in the single vortex state $L_{z}=N$ follows the GOE distribution, as shown in Fig. 6(c). At large value of $L$, it saturates to a value $\Delta_{\infty}=0.187,0.300$ and 0.4890.4890.489 for $N=12,16,20$, respectively. These values are smaller than in the case of moderate interaction case $\Delta_{\infty}=0.916,0.403,0.657$ for the same number of bosons. This shows that the rotating system with strong interaction is strongly chaotic.
@@
We also present the gap ratio distribution $P(r)$ and the spectral average $\Delta_{3}(L)$ for strong interaction regime ($g_{2}=7.301$) corresponding to angular momentum $L_{z}=2N$ and $L_{z}=3N$, as shown in Figs. 7 and 8, respectively. In Fig. 7, we observe that the short-range correlation as measured by $P(r)$ follows the GOE distribution for both $L_{z}=2N$ (upper panel) and $L_{z}=3N$ (lower panel). Further, it is observed from Fig. 8 that $\Delta_{3}(L)$ as a measure of the long-range correlation exhibits good agreement with the GOE distribution for small values of energy interval $L$ both for $L_{z}=2N$ (left) and $L_{z}=3N$ (right). Thus, the analysis of both short-range and long-range correlations exhibits signatures of quantum chaos for $L_{z}=2N,3N$. This may be attributed to the interplay between strong two-body interaction, rotation and the number of bosons.
@@ V Conclusion
Thus, the interplay between the trap energy and interaction energy governs the transition of the system from regular to chaotic behavior, while rotation further amplifies the degree of chaos. Our study on trapped interacting bosons signifies the universal applicability of RMT in describing the spectral correlations in quantum many-body systems. We believe that these findings will contribute to advancing the understanding of quantum chaos in ultracold Bose systems.
The present work may be readily extended to explore the spectral form factor (SFF) to investigate the long-time behavior of the trapped interacting Bose system.
@@ Acknowledgements.
Mohd Talib thanks the Ministry of Social Justice and Empowerment, Government of India, for providing the Junior Research Fellowship (JRF) to carry out this research work (NBCFDC Ref. No.: 221610104398).

### v2

@@ IV.2 Rotating case
Figure 5: (Color online) Upper panel: Nearest-neighbor spacing distribution $P(s)$; Lower panel: Distribution of the ratio of consecutive level spacings $P(r)$ for the strong interaction regime with $g_{2}=3.669$ in the single-vortex state $L_{z}=N$ for $N=12$, $16$, and $20$. The histogram in each graph represents our numerical result for the lowest 100 energy levels. The blue line shows Poisson, the orange dashed line GOE, and the green dash-dotted line the Brody distribution with fitted parameter $b$.
(a) $L_{z}=N$, $g_{2}=0.3669$
(b) $L_{z}=N$, $g_{2}=0.3669$
(c) $L_{z}=N$, $g_{2}=3.669$
(d) $L_{z}=N$, $g_{2}=3.669$
Figure 6: (Color online) Spectral average $\langle\Delta_{3}(L)\rangle$ vs. $L$ and number variance $\Sigma^{2}(L)$ vs. $L$ for moderate interaction with $g_{2}=0.3669$ (upper panel) and strong interaction with $g_{2}=3.669$ (lower panel) regimes in the single-vortex state $L_{z}=N$ for $N=12$, $16$, and $20$. Blue circles, green squares, and red diamonds show numerical results. Black dashed and magenta dash-dotted lines correspond to Poisson and GOE results, respectively.
Table 6: The mean gap ratio with estimated error for the two subsets of energy levels—the lower 1–50 and the higher 51–100 levels—is calculated for different number of bosons in the rotating single-vortex case ($L_{z}=N$) in both moderate ($g_{2}=0.3669$) and strong ($g_{2}=3.669$) interaction regimes.
Number of bosons, NN Number of levels Mean gap ratio $\langle r\rangle$
$g_{2}=0.3669$ $g_{2}=3.669$
12 1–50 0.518 $\pm$ 0.041 0.569 $\pm$ 0.035
51–100 0.478 $\pm$ 0.038 0.545 $\pm$ 0.038
16 1–50 0.529 $\pm$ 0.039 0.520 $\pm$ 0.037
51–100 0.472 $\pm$ 0.040 0.486 $\pm$ 0.041
20 1–50 0.473 $\pm$ 0.035 0.564 $\pm$ 0.035
51–100 0.517 $\pm$ 0.036 0.545 $\pm$ 0.030
Poisson 0.386
GOE 0.530
@@
Figure 7: (Color online) The distribution of the ratio of consecutive level spacings $P(r)$ for strong interaction regime with $g_{2}=3.669$ for $N=12,16,20$ in $L_{z}=2N$ (Upper Panel) and $L_{z}=3N$ (Lower Panel). The histogram in each graphs represents our numerical result for the lowest 100 energy levels. The blue line corresponds to the Poisson distribution, the orange dashed curve to the GOE distribution.
(a) $L_{z}=2N$, $g_{2}=3.669$
(b) $L_{z}=3N$, $g_{2}=3.669$
Figure 8: (Color online) The spectral average $\langle\Delta_{3}(L)\rangle$ vs $L$ is presented for strong interaction regime with $g_{2}=3.669$ for different number of bosons $N=12,16$ and $20$ in $L_{z}=2N$ (Left) and $L_{z}=3N$ (Right). The blue circle, green square and red diamond lines are our numerical results of $\langle\Delta_{3}(L)\rangle$ for $N=12,16$ and $20$, respectively, for the lowest 100 energy levels . For reference, we have also drawn $\langle\Delta_{3}(L)\rangle$ as a function of $L$, corresponding to the Poisson distribution (black dashed line) and the GOE distribution (magenta dash-dot line).
IV.2.1 Short-range correlations
The distributions $P(s)$ and $P(r)$ for various values of $L_{z}=N$ corresponding to number of bosons $N=12,16$ and $20$, are presented in Fig. 4 for the moderate interaction regime and in Fig. 5 for the strong interaction regime.
Moderate interaction regime
For number of bosons $N=12,16$, 2020 and for the single vortex state $L_{z}=N$, the NNSD $P(s)$ distribution exhibits GOE behavior with Brody parameter $b=0.52,0.64$ and $0.64$, respectively, as shown in Figs. 4(a)-4(c). This indicates that these systems exist in a weakly chaotic regime, retaining substantial degree of regularity.
The gap ratio distribution $P(r)$ for $N=12,16$ and $20$ in the single vortex state $L_{z}=N$ are presented in Figs. 4(d)-4(f) with mean gap ratio $\langle r\rangle=0.498,0.494$ and $0.496$, respectively, indicating that these systems remain in a weakly chaotic regime. The mean gap ratio corresponding to the lower (115050) and the higher (5151100100) energy levels of the spectrum is listed in the fourth column of Table 6. The lower part of the spectrum exhibits chaotic behavior consistent with GOE distribution ($\langle r\rangle\approx 0.530$), while the higher energy levels display weaker chaoticity, for $N=12$ and $16$. However, for $N=20$, the upper part (51-100) of the spectrum exhibits chaotic behavior consistent with GOE, while the lower part (1-50) is two standard deviations below GOE.
@@ IV.2.1 Short-range correlations
The gap ratio distribution $P(r)$ for $N=12,16,20$ in the single vortex state $L_{z}=N$ follows the GOE distribution with mean gap ratio $\langle r\rangle=0.555,0.505$ and $0.549$, respectively, as shown in Figs. 5(d)-5(f). This indicate that these systems show strong signature of quantum chaos.
The Brody parameter $b$ and the mean gap ratio $\langle r\rangle$ values for each of the number of bosons $N$ in the single vortex state $L_{z}=N$ with moderate and strong interaction regimes are listed in Table 3 and 4. The mean gap ratio for the lower (115050) and higher (5151100100) subsets of energy levels is presented in the fifth column of Table 6. Both the lower and the higher parts of the spectrum exhibit chaotic behavior consistent with the GOE distribution ($\langle r\rangle\approx 0.530$).
IV.2.2 Long-range correlations
Moderate interaction regime
The spectral average $\Delta_{3}(L)$ for $N=12,16,20$ in the single vortex state $L_{z}=N$ is shown in Fig. 6(a). It is observed that for small energy intervals $L$, the $\Delta_{3}(L)$ statistic exhibits GOE like behavior, while for large $L$, it approaches a saturation value. The saturated values for $N=12,16$ and $20$ are found to be $\Delta_{\infty}$= $0.916$, $0.403$ and $0.657$, respectively. The value of $\Delta_{\infty}$ is modulated by the number of bosons $N$ in the moderate interaction regime which may be attributed to the interplay between the number of bosons and the interaction strength.
Strong interaction regime
For small values of energy interval $L$, the numerical data of $\Delta_{3}(L)$ statistic for $N=12,16,20$ in the single vortex state $L_{z}=N$ follows the GOE distribution, as shown in Fig. 6(c). At large value of $L$, it saturates to a value $\Delta_{\infty}=0.187,0.300$ and $0.489$ for $N=12,16,20$, respectively. These values are smaller than in the case of moderate interaction case $\Delta_{\infty}=0.916,0.403,0.657$ for the same number of bosons. This shows that the rotating system with strong interaction is strongly chaotic.
@@ IV.2.2 Long-range correlations
We also present the gap ratio distribution $P(r)$ and the spectral average $\Delta_{3}(L)$ for strong interaction regime ($g_{2}=3.669$) corresponding to angular momentum $L_{z}=2N$ and $L_{z}=3N$, as shown in Figs. 7 and 8, respectively. In Fig. 7, we observe that the short-range correlation as measured by $P(r)$ follows the GOE distribution for both $L_{z}=2N$ (upper panel) and $L_{z}=3N$ (lower panel). Further, it is observed from Fig. 8 that $\Delta_{3}(L)$ as a measure of the long-range correlation exhibits good agreement with the GOE distribution both for $L_{z}=2N$ (left) and $L_{z}=3N$ (right). Thus, the analysis of both short-range and long-range correlations exhibits signatures of quantum chaos for $L_{z}=2N,3N$. This may be attributed to the interplay between strong two-body interaction, rotation and the number of bosons.
@@ V Conclusion
Thus, the interplay between the interaction energy and the trap energy governs the transition of the system from regular to chaotic behavior, while rotation further amplifies the degree of chaos. Our study on trapped interacting bosons signifies the universal applicability of RMT in describing the spectral correlations in quantum many-body systems. We believe that these findings will contribute to advance the understanding of quantum chaos in ultracold Bose systems.
The present work may be readily extended to explore the spectral form factor to investigate the late time behavior of the trapped interacting Bose system. It would also be worthwhile to investigate how the energy-level statistics of an extended Bose-Hubbard model [67] evolve as the strength of the cavity-induced infinite-range interactions is increased.
@@ Acknowledgements.
Mohd Talib thanks the Ministry of Social Justice and Empowerment, Government of India, for providing the Junior Research Fellowship (JRF) (NBCFDC Ref. No.: 221610104398).